OpenStudy (anonymous):

You buy 2 bags of chips and 3 boxes of popcorn for $5.13. You then buy another bag of chips and 2 more boxes of popcorn for $3.09. What is the cost of a bag of chips and a box of popcorn?

OpenStudy (anonymous):

let x represents cost of one bag of chips nd y represents cost of one box... then 2x +3y=5.13 nd x+2y=3.09.. solve for x nd y

OpenStudy (anonymous):

????

OpenStudy (anonymous):

just solve for it and that gives me my answer

OpenStudy (anonymous):

u got two equations with 2 variables.. find x nd y frm here by subsitution method... this will give u ur answer

OpenStudy (anonymous):

yes @Ashley_Cute

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

y=5.13

OpenStudy (anonymous):

Solve x+2y=3.09 for x and then plug the result into the first equation which leaves one equation in one unknown.

OpenStudy (anonymous):

no @Ashley_Cute

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

i got x+y=1.545

OpenStudy (anonymous):

how i get rid of y

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

x +y =1.545.. so x= 1.545-y.. put this value of x in any equation.. find y

OpenStudy (anonymous):

Move 2y to the RHS of the second equation. Do you get: x=3.09-2y ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now replace each x in the first equation with "3.09-2y"

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

while im solving this do i combine like terms the y's

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

Yes, combine all y terms.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

99cents

OpenStudy (anonymous):

Yes, x = 99 cents Replace x in the second equation with 0.99 and then solve for y.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

1.05

OpenStudy (anonymous):

Yes, very good. When you have two unknowns, like x and y, then it takes two equations to find x and y.

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

np and thank you